OpenStudy (anonymous):

How would you write 1.9398 x 10^4 in scientific notation?

OpenStudy (anonymous):

thats in scientific notation already

OpenStudy (anonymous):

Yea, that's what I thought too, but it's telling me it's not the right answer.

OpenStudy (anonymous):

its defined as N.N... x 10^n where the first number is between 1 and 10 have you got the full question?

OpenStudy (anonymous):

Yes would you like me to give you the complete question?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

The mass of Neptune is about 1.03 x 10^23 metric tons. The mass of the sun is about 1.998 x 10^27 metric tons. About how many times the mass of Neptune is the mass of the sun? My answer is 1.9398 x 10^4 which I think is the right answer but when I input it, it says incorrect and to divide carefully and it must be greater than one but less than 10. Doesn't make sense, unless their is a glitch in the system.

OpenStudy (anonymous):

try 1.94x10^4

OpenStudy (anonymous):

1.998 / 1.03 = 1.9398 to 4 decimal places and 10^27 / 10^23 = 10^4 - correct did it say to how many decimal places it wanted? - but as you say, there might be a glitch in the system

OpenStudy (anonymous):

It says to round to the nearest thousandth

OpenStudy (anonymous):

right well thats 1.940

OpenStudy (anonymous):

OMG, DUH

OpenStudy (anonymous):

it it's nearest thousandth try 1.940x 10^4

OpenStudy (anonymous):

How stupid of me! Thank you

OpenStudy (anonymous):

no probs